# Millimeter-wave Radio SLAM: End-to-End Processing Methods and Experimental Validation

### v1

@@ VII Conclusions
In this article, we addressed the timely notion of mmWave radio SLAM from end-to-end processing perspective. We first proposed a novel SVD-based estimation approach for acquiring the AoAs and AoDs of the involved propagation paths. The method operates on BRSRP measurements and does not need information of the antenna system beampatterns or the involved beamformers, while offering built-in robustness against antenna sidelobes. Secondly, a new snapshot SLAM method was also proposed, to jointly estimate the locations of the landmarks and the UE, offering improved robustness and identifiability compared to prior-art. The performance of the proposed methods was comprehensively assessed through ray tracing and true measurement data at 60 GHz. The results show that the methods outperform the relevant prior-art, with the end-to-end performance being comparable or even better compared to sequential filtering solutions while offering substantially reduced complexity. Finally, we provide the measured 60 GHz data openly available for the research community. Our future work considers extending the proposed angle estimation methods to include multi-bounce detection and estimation with rank-1 singular matrices, thus facilitating further evolved SLAM solutions beyond the common single-bounce approaches.

### v2

@@ VII Conclusions
In this article, we addressed the timely notion of mmWave radio SLAM from end-to-end processing perspective, under minimal knowledge regarding the array parameters and user mobility. We first proposed a novel SVD-based estimation approach for acquiring the AoAs and AoDs of the involved propagation paths. The method operates on BRSRP measurements and does not need information on the complex antenna patterns or the underlying steering vectors and beamforming weights, while offering built-in robustness against antenna sidelobes. The method relies only on known physical correspondence between the beam indices and the beamforming angles, and the sparsity of the beamformed mmWave channel. Secondly, a new snapshot SLAM method was also proposed, to jointly estimate the locations of the landmarks and the UE, offering improved robustness and identifiability compared to prior-art. The performance of the proposed methods was comprehensively assessed through ray tracing and true measurement data at 60 GHz. The results show that the methods outperform the relevant prior-art, with the end-to-end performance being comparable or even better compared to sequential filtering solutions while offering substantially reduced complexity. Finally, we provide the measured 60 GHz data openly available for the research community. Our future work considers extending the proposed angle estimation methods to include multi-bounce detection and estimation with rank-1 singular matrices, thus facilitating further evolved SLAM solutions beyond the common single-bounce approaches.
Biographies
[Uncaptioned image] Elizaveta Rastorgueva-Foi is a doctoral candidate at the Unit of Electrical Engineering at Tampere University, Finland, where she also received her M.Sc. degree in electrical engineering in 2019. Her research interests include statistical signal processing, positioning and location-aware communications in mobile networks with an emphasis on 5G and beyond systems.
[Uncaptioned image] Ossi Kaltiokallio (M’23) received the Ph.D. degree in Electrical Engineering in 2017 from Aalto University, Finland. He has held postdoctoral fellow position at the University of Utah, Aalto University and Tampere University. Currently he is a Senior Research Fellow with the Electrical Engineering Unit, Tampere University, Finland. His current research interests lie at the intersection of statistical signal processing and wireless networking for improving radio-based positioning and sensing technologies.
[Uncaptioned image] Yu Ge (S’20) received his B.E. degree from Zhejiang University, Hangzhou, China, in 2017, and M.Sc. degree from the KTH Royal Institute of Technology, Stockholm, Sweden, in 2019. He is currently a Ph.D. candidate in the Department of Electrical and Engineering at Chalmers University of Technology, Sweden. His research interests include integrated communication and sensing, wireless positioning systems, simultaneous localization and mapping, and multi-object tracking, particularly in 5G and Beyond 5G scenarios.
[Uncaptioned image] Matias Turunen is a researcher and laboratory specialist at the Department of Electrical Engineering, Tampere University (TAU), Finland. His research interests include inband full-duplex radios with an emphasis on analog RF cancellation, OFDM radar, and 5G New Radio systems.
[Uncaptioned image] Jukka Talvitie (S’09, M’17) received the M.Sc. and D.Sc. degrees from the Tampere University of Technology, Finland, in 2008 and 2016, respectively. He is currently a University Lecturer with the Unit of Electrical Engineering, Tampere University (TAU), Finland. His research interests include signal processing for wireless communications, radio-based positioning and sensing, radio link waveform design, and radio system design, particularly concerning 5G and beyond mobile technologies.
[Uncaptioned image] Bo Tan received his PhD from the Institute for Digital Communications (IDCOM), The University of Edinburgh, UK, in Nov. 2013. From 2012 to 2016, he conducted multiple postdoctoral research projects at the University College London and the University of Bristol, UK, contributing to passive radar design and applications in security and healthcare. From 2017 to 2018, he was a lecturer at Coventry University, UK. Since 2023, he has been a tenure-track associate professor at Tampere University, Finland. His research interests include radio sensing and connectivity for intelligent machines.
[Uncaptioned image] Musa Furkan Keskin received the B.S., M.S., and Ph.D. degrees from the Department of Electrical and Electronics Engineering, Bilkent University, Ankara, Turkey, in 2010, 2012, and 2018, respectively. He is currently Research Specialist with the department of Electrical Engineering at Chalmers University of Technology, Gothenburg, Sweden. His current research interests include integrated sensing and communications, RIS-aided localization and sensing, and hardware impairments in beyond 5G/6G systems.
[Uncaptioned image] Henk Wymeersch (S’01, M’05, SM’19, F’24) obtained the Ph.D. degree in Electrical Engineering/Applied Sciences in 2005 from Ghent University, Belgium. He is currently a Professor of Communication Systems with the Department of Electrical Engineering at Chalmers University of Technology, Sweden. He is currently a Senior Member of the IEEE Signal Processing Magazine Editorial Board. During 2019-2021, he was an IEEE Distinguished Lecturer with the Vehicular Technology Society. His current research interests include the convergence of communication and sensing, in a 5G and Beyond 5G context.
[Uncaptioned image] Mikko Valkama [S’00, M’01, SM’15, F’22] received his M.Sc. (Tech.) and D.Sc. (Tech.) degrees (both with honors) from Tampere University of Technology, Finland, in 2000 and 2001, respectively. In 2003, he was with the Communications Systems and Signal Processing Institute at SDSU, San Diego, CA, as a visiting research fellow. Currently, he is a Full Professor and the Head of the Unit of Electrical Engineering at the newly formed Tampere University, Finland. His general research interests include radio communications, radio localization, and radio-based sensing, with particular emphasis on 5G and 6G mobile radio networks.